# Accelerated Mirror Descent for Non-Euclidean Star-convex Functions

### v1

@@ Abstract
Acceleration for non-convex functions has been an important problem in optimisation. We revisit star-convex functions, which are strictly unimodal on all lines through a minimizer. In [18], the authors accelerate gradient descent for star-convex functions with gradients that are Lipschitz with respect to the Euclidean norm in an unconstrained domain. In this paper, we introduce a new assumption about the regularity of the derivative of a general norm and we accelerate mirror descent for this class of normed spaces. We show that, under it, our algorithms show sharp convergence rates for star-convex functions with $\alpha$-Hölder continuous gradients. We also prove that our convergence rate is near optimal for $p$-norms.
@@ 1 Introduction
Nesterov’s accelerated gradient descent [26] has proven to be a powerful tool for first-order optimization, significantly improving algorithmic performance. Nowadays several extensions of accelerated methods exist including coordinate descent methods [27, 12], distributed gradient descent [30] and proximal methods [13] with applications across widespread domains such as image deblurring [2], compressed sensing [3] and deep learning [31].
@@
Theoretical results analysing accelerated methods mainly focus on smooth convex optimisation, demonstrating optimal convergence and improved performance against traditional gradient methods, see [26] and [6] for further references. However, this framework becomes less applicable to modern machine learning applications due to their typically non-convex energy landscapes. Attempts to understand acceleration for non-convex functions has been made, for instance, by leveraging the Łojasiewicz inequalities [5]. Further classes of non-convex functions are discussed in [7] and [18](Appendix A.1).
@@
[28] relaxes the convexity assumption to so-called star-convex functions, a class of "structured" non-convex function, which are strictly unimodular on all lines through a minimizer. As convex functions exhibit unimodularity along all lines, star-convexity represents a weaker condition. [23] provide examples of functions that are star-convex but not convex. Additionally, [36] and [22] provide evidence that neural network loss functions often satisfy the star-convexity (but not the convexity) assumption, emphasizing its relevance for modern machine learning applications.
@@
For star-convex functions, [18] present an accelerated framework using binary search techniques that is nearly optimal provided that the objective function has a Lipschitz gradient in Euclidean norm. However, recent machine learning results such as [32, 1] minimize the objective functions in non-Euclidean norms, which are not covered by the findings in [18]. This gap inspires us to extend and generalize their algorithm for arbitrary normed spaces. We identify a new regularity condition for norms (Definition 2.7) that enables acceleration. In regular normed spaces, we combine a generalized "line-search" style algorithm [18] (Algo 2) with accelerated mirror descent to minimize star-convex functions with $\alpha$-Hölder continuous gradients (also known as weakly smooth functions, Definition 2.4). This indicates that compensating for the non-convexity in the objective function requires regularity in the norm rather than the function itself. Our finding challenges the common belief that the properties of the objective function are more crucial than those of the norm.
We demonstrate that in case of $p$-norms, which are regular for $p>1$ (see Lemma 2.1), our algorithm is nearly optimal. By showing that non-convex functions can be solved by accelerated algorithms in non-Euclidean geometry, we extend the traditional understanding of which functions admit acceleration. Our work suggests that more non-convex functions can be solved efficiently, potentially sparking increased interest in this area of research.
@@ 1.1 Literature review
Previous results on star-convex functions primarily focus on non-accelerated algorithms. [14] and [17] show convergence guarantees for stochastic gradient descent, [28] analyse cubic regularization and [23] investigate a cutting plane method. In [20], the star-convexity assumption is applied to stochastic and online learning settings, showing regret bounds for first-order methods, leading to a $\mathcal{O}(1/T)$ convergence rate (with $T$ being the number of gradient steps) in case that either the objective function or its gradient is Lipschitz.
Acceleration on star-convex functions has been studied in, e.g., [15] for low-dimensional subspace optimization and in [35] for an one-dimensional line search-type of algorithm, both showing convergence rates to an $\epsilon$-optimal point. However, as noted in [18], these algorithms can be limited by their potentially high computational costs for minimizing general star-convex functions.
[18] proposes an accelerated framework using binary search, showing near-optimal performance, provided the $\tau$-star-convex functions (see Definition 2.1) are $L$-smooth (see Definition 2.3) with respect to the Euclidean norm in a domain bounded by $R$. Similar to [15], their algorithm finds an $\epsilon$-optimal solution in $\mathcal{O}(\epsilon^{-1/2}L^{1/2}\tau R\log(LR^{2}/\epsilon))$ iterations. For smooth star-convex functions, they also provide a lower bound of $\Omega(L^{1/2}\tau R\epsilon^{-1/2})$.
While the accelerated framework in [18] is limited to smooth star-convex functions with respect to the Euclidean norm, Nesterov’s accelerated framework ([11, 24]) applies to general normed spaces under a weak smoothness (see Definition 2.4) assumption.
@@ 1.2 Our contributions
In this paper, we generalize the accelerated framework introduced in [18] to encompass weakly smooth star-convex functions with respect to a general norm. Addressing the challenges arising from the non-convexity of the objective function and the non-Euclidean geometry of the norm, we introduce new concepts, such as regular norms, and develop novel proof strategies for the algorithms convergence. Our key contributions are summarized as follows:
Acceleration on regular norms. To extend the acceleration framework of [18] from Euclidean norms to general norms, we identify sufficient conditions on the derivative of the squared norm $\|\cdot\|^{2}$. Introducing the notion of regular norms, we demonstrate that star-convexity and weak smoothness suffice to apply our generalized accelerated framework. Our results underscore the critical role of norm regularity, rather than the function properties, in achieving acceleration, encouraging further research in this direction.
Extended binary search. We generalize the binary search algorithm introduced in [18], originally designed for smooth functions to encompass weakly smooth objective functions. Additionally, we present a streamlined convergence proof (Thm 4.1) for our generalized algorithm.
Nearly optimal performance for $p$-norm. For $p$-norms, we extend the convergence rates in [24] from convex to star-convex functions. According to the lower bounds in [16], our rates are nearly optimal, up to some logarithmic factor. Table 1 categorises our results in the existing literature.
@@
$(1,2]$ Upper $\mathcal{O}_{\kappa}\left(\left(\frac{LR^{\kappa}}{\epsilon}\right)^{\frac{2}{3\kappa-2}}\right)$ $\mathcal{O}_{\kappa}\left(\left(\frac{L\tau^{2}R^{\kappa}}{\epsilon}\right)^{\frac{2}{3\kappa-2}}\log\Big{(}\frac{LR^{\kappa}}{\epsilon}\Big{)}\right)$
[24] Our work (Corollary 3.3)
Lower $\Omega_{\kappa}\left(\left(\frac{LR^{\kappa}}{\epsilon[\ln d]^{\kappa-1}}\right)^{\frac{2}{3\kappa-2}}\right)$
[16]
$[2,\infty]$ Upper $\mathcal{O}_{p,\kappa}\left(\left(\frac{LR^{\kappa}}{\epsilon}\right)^{\frac{p}{\kappa p+\kappa-p}}\right)$ $\mathcal{O}_{p,\kappa}\left(\left(\frac{L\tau^{p}R^{\kappa}}{\epsilon}\right)^{\frac{p}{\kappa p+\kappa-p}}\log\Big{(}\frac{LR^{\kappa}}{\epsilon}\Big{)}\right)$
[24] Our work (Corollary 3.4)
Lower $\Omega_{p,\kappa}\left(\left(\frac{LR^{\kappa}}{\epsilon}\right)^{\frac{p}{\kappa p+\kappa-p}}\right)$
[16]
Table 1: Summary of the relation between related existing work and our work for $\tau$-star-convex functions. $R$ is the radius of the domain of the $\tau$-star-convex objective function, $(L,\kappa)$ is its smoothness with respect to a $p$-norm $\|\cdot\|_{p}$ with $p>1$ and $\epsilon$ the accuracy of the algorithm. Since convex functions are $1$-star-convex, the lower bounds from [16] apply to both.
Paper outline. In Section 2, we outline the problem setting and provide preliminaries, including the norm regularity assumption and its properties. Section 3 presents our generalized accelerated algorithm and discusses its convergence guarantees for various settings. Section 4 introduces our binary search algorithm and its running time, and Section 5 concludes with a discussion on future research directions. All proofs are provided in the Appendix.
@@ 2.1 Definitions
In this paper, we consider the class of $\tau$-star-convex functions, as introduced in [17], which are defined as follows.
@@
A continuously differentiable function $F:\mathcal{X}\xrightarrow{}\mathbb{R}$ is said to be $\tau$-star-convex for $\tau>0$ if there exists $x_{\star}\in\mathcal{X}$ such that for all $x\in\mathcal{X}$
$\tau\langle\nabla F(x),x-x_{\star}\rangle\geq F(x)-F(x_{\star}).$ (1)
Note that, all convex functions are $1$-star-convex [20]. If included, 00-star convexity is trivial.
@@
For a continuously differentiable function $F:\mathcal{X}\xrightarrow{}\mathbb{R}$, the Bregman divergence is defined as
@@
Definition 2.3 (Smoothness).
[…]
$|D_{F}(x,y)|\leq\frac{L}{2}\|x-y\|^{2}.$
@@
When $\kappa=2$, weak smoothness becomes the usual smoothness. In the general case, we would refer to $\kappa<2$.
To develop our mirror descent algorithm, we use a distance-generating function (see Assumption 1), that fulfills strong or uniform convexity assumptions with respect to a norm, defined as follows.
Definition 2.5 (Strong convexity).
[…]
$D_{F}(x,y)\geq\frac{\mu}{2}\|x-y\|^{2}.$
@@
When $q=2$, it is equivalent to the usual strong convexity. In the general case, we would have $q>2$.
2.2 Norm regularity assumption
Without convexity of the objective function, acceleration is possible with additional assumptions, usually by constraining the class of objective functions. In our paper, we introduce a new regularity property for norms that we will show is a sufficient condition for acceleration.
[…]
$\phi(x):=\nabla\left(\frac{\|x\|^{2}}{2}\right).$
Definition 2.7 (Regular norm).
[…]
$\begin{cases}\langle\phi(x),x\rangle&\geq\|x\|^{2}\\ \|\phi(x)\|_{\ast}&\leq C\|x\|.\end{cases}$
Given that our setting is finite dimensional, the equivalence of all norms suggests a potentially large family of regular norms. However, to provide dimension-independent performance guarantees for our algorithm, we require the norms to be $C$-regular, with $C>0$ being independent of the dimension.
Since $\phi$ is a homogeneous function by definition, the regularity can be rewritten as
$\begin{cases}\inf_{y\in\mathbb{U}}\quad\langle\phi(y),y\rangle\geq 1\\ \sup_{y\in\mathbb{U}}\quad\|\phi(y)\|_{\ast}\leq C,\\ \end{cases}$
where $\mathbb{U}$ denotes the unit sphere. The first condition ensures compatibility of the norm with the scalar product, while the second condition requires the norm’s derivative operator to be bounded. These two properties are crucial in our accelerated framework (Algorithm 1), especially for weakly smooth objective functions. They allow for the choice of aggressive step-sizes and the derivation of accelerated convergence rates.
The following lemmas provide examples of norms that are $C$-regular with constant $C>0$ not depending on the input dimension $d$. The proofs can be found in Appendix A.
Lemma 2.1.
For $p>1$, the $p$-norm $\|\cdot\|_{p}$ is $1$-regular. The matrix Schatten $p$-norm is also $1$-regular.
Another example of a regular norm is the following: Fix $\lambda\in[0,1]$, and consider a norm $\|\cdot\|_{a}$ defined on $\mathbb{R}^{d_{a}}$ and a norm $\|\cdot\|_{b}$ defined on $\mathbb{R}^{d_{b}}$. The composite norm $\|\cdot\|_{a\circ b}$ is defined on $\mathbb{R}^{d_{a}+d_{b}}$ by
$\|x\|_{a\circ b}=\left\|(x_{a},x_{b})^{T}\right\|_{a\circ b}:=\sqrt{\lambda\|x_{a}\|_{a}^{2}+(1-\lambda)\|x_{b}\|_{b}^{2}},$
where $x_{a}:=x_{1:d_{a}}$ and $x_{b}:=x_{(d_{a}+1):(d_{a}+d_{b})}$.
Lemma 2.2.
Let norm $\|\cdot\|_{a}$ be $C_{a}$-regular and norm $\|\cdot\|_{b}$ be $C_{b}$-regular, then the composite norm $\|\cdot\|_{a\circ b}$ is $\max\{C_{a},C_{b}\}$-regular.
Lemma 2.2 indicates that our algorithm provides convergence guarantees not only for $p$-norms, but also for compositions of regular norms. Combined with Lemma 2.1, this shows that our algorithm is applicable for a wide range of non-Euclidean norms. For instance, since $\|\cdot\|_{2}$ and $\|\cdot\|_{1.5}$ are $1$-regular according to Lemma 2.1, Lemma 2.2 implies that the following composite norm is is also $1$-regular
$\|x\|_{2\circ 1.5}:=\sqrt{0.5\|x_{1:d/2}\|^{2}_{2}+0.5\|x_{(d/2+1):d}\|^{2}_{1.5}}.$ (2)
@@ 3 Main results
In this section, we develop our accelerated optimization algorithm for weakly smooth star-convex functions. We start by summarizing our assumptions.

### v2

@@ Abstract
Acceleration for non-convex functions is a fundamental challenge in optimisation. We revisit star-convex functions, which are strictly unimodal on all lines through a minimizer. [pmlr-v125-hinder20a] accelerate unconstrained star-convex minimization of functions that are smooth with respect to the Euclidean norm. To do so, they add a certain binary search step to gradient descent. In this paper, we accelerate unconstrained star-convex minimization of functions that are weakly smooth with respect to an arbitrary norm. We add a binary search step to mirror descent, generalize the approach and refine its complexity analysis. We prove that our algorithms have sharp convergence rates for star-convex functions with $\alpha$-Hölder continuous gradients and demonstrate that our rates are nearly optimal for $p$-norms.
@@ 1 Introduction
Accelerated gradient descent by [Nesterov1983AMF] has proven to be a powerful tool for first-order optimization, significantly improving algorithmic performance. Nowadays several extensions of accelerated methods exist including coordinate descent methods [coord1, coord2], distributed gradient descent [Quli16] and proximal methods [prox1] with applications across widespread domains such as image deblurring [BT09], compressed sensing [Becker11] and deep learning [acc-ex2].
@@
Theoretical results analysing accelerated methods mainly focus on smooth convex optimisation, demonstrating optimal convergence and improved performance against traditional gradient methods, see [Nesterov1983AMF],  [bubeck2014convex] and [AGJ18] for further references. However, these assumptions are violated in modern machine learning applications due to their non-convex energy landscapes. Attempts to understand acceleration for non-convex functions are made, for instance, by leveraging the Łojasiewicz inequalities [bolte10]. Further classes of non-convex functions are discussed by [Carmon2017ConvexUP] and [pmlr-v125-hinder20a].
@@
[Nesterov2006CubicRO] relax the convexity assumption to so-called star-convex functions, a class of structured non-convex functions that are strictly unimodal on all lines through a minimizer. While convex functions exhibit unimodality along all lines, star-convexity is a weaker condition. [Lee16] provide examples of functions that are star-convex but not convex. Additionally, [Zhou2019SGDCT] and [pmlr-v80-kleinberg18a] present evidence that neural network loss functions often are star-convex (but not convex), emphasizing its relevance for modern machine learning applications.
@@
For star-convex functions, [pmlr-v125-hinder20a] present an accelerated framework using binary search techniques. Their method is nearly optimal for objective functions having a Lipschitz gradient with respect to the Euclidean norm. However, recent machine learning results such as by [Yang2015, Adil19] minimize the objective functions in non-Euclidean norms, which are not covered by the findings of [pmlr-v125-hinder20a]. This gap inspires us to extend and generalize their algorithm to arbitrary normed spaces. We combine a generalized “line-search” style algorithm [pmlr-v125-hinder20a] with accelerated mirror descent to minimize star-convex functions with $\alpha$-Hölder continuous gradients (also known as weakly smooth functions, see Definition 2.3). We demonstrate that the general non-Euclidean acceleration framework is not related to the convexity of the objective function, as previously theorized. Rather, a key factor appears to be the inherent regularity characteristic of norms (Fact A.1). Our finding challenges the common belief that the properties of the objective function are more crucial than those of the norm.
We demonstrate that in case of $p$-norms, our algorithm is nearly optimal for $p>1$. Our approach further extends to the $L_{1}$ norm though with an additional dimensionality-dependent cost tied to challenges in constructing an appropriate so-called distance generating function. By showing that non-convex functions can be minimized using accelerated algorithms in non-Euclidean geometry, we broaden the conventional understanding of which functions admit acceleration. Our findings suggest that a wider class of non-convex functions can be efficiently minimized, potentially driving further research in this direction.
@@ 1.1 Literature review
Previous results on star-convex functions primarily focus on non-accelerated algorithms. [Gower2020SGDFS] and [Hardt2016GradientDL] show convergence guarantees for stochastic gradient descent, [Nesterov2006CubicRO] analyze cubic regularization and [Lee16] investigate a cutting plane method. [joulani17] apply the star-convexity assumption to stochastic and online learning settings, showing regret bounds for first-order methods, leading to a $\mathcal{O}(1/T)$ convergence rate (with $T$ being the number of gradient steps) in case that either the objective function or its gradient are Lipschitz.
Acceleration on star-convex functions has been studied by, e.g., [Guminov2018AcceleratedMF] for low-dimensional subspace optimization and by [Nest19] for an one-dimensional line search-type of algorithm, both showing convergence rates to an $\epsilon$-optimal point. However, as noted by [pmlr-v125-hinder20a], these algorithms can be limited by their potentially high computational costs for minimising general star-convex functions.
[pmlr-v125-hinder20a] propose an accelerated framework using binary search, showing near-optimal performance for $\tau$-star-convex functions (see Definition 2.1) that are $L$-smooth (see Definition 2.3) with respect to the Euclidean norm in a domain bounded by radius $R$. Similar to [Guminov2018AcceleratedMF], their algorithm finds an $\epsilon$-optimal solution in $\mathcal{O}(\epsilon^{-1/2}L^{1/2}\tau R\log(LR^{2}/\epsilon))$ iterations. For smooth star-convex functions, they also provide a lower bound of $\Omega(L^{1/2}\tau R\epsilon^{-1/2})$.
While the accelerated framework of [pmlr-v125-hinder20a] is limited to star-convex functions that are smooth with respect to the Euclidean norm, Nesterov’s accelerated framework [Elster1993ModernMM, NEMIROVSKII198521] for convex functions applies to general normed spaces under a weaker smoothness assumption (Definition 2.3 below).
@@ 1.2 Our contributions
In this paper, we generalize the accelerated framework introduced by [pmlr-v125-hinder20a] to encompass weakly smooth star-convex functions with respect to a general norm. To address the challenges arising from the non-convexity of the objective function and the non-Euclidean geometry of the norm, we present a new convergence proof by using a regularity property of norms, involving novel arguments for bounded iterates to quantify convergence. Our key contributions are summarized as follows:
Acceleration for arbitrary norms. To extend the acceleration framework of [pmlr-v125-hinder20a] beyond Euclidean norms to general norms, we examine regularity properties of the derivative of the squared norm $\|\cdot\|^{2}$. We show that star-convexity and weak smoothness suffice to apply our generalized accelerated framework. Our results underscore that the key to acceleration lies in the convexity and homogeneity of norms, rather than convexity of the objective function.
Extended binary search. We generalize the binary search algorithm introduced by [pmlr-v125-hinder20a], originally designed for functions that are smooth with respect to the Euclidean norms, to encompass objective functions that are weakly smooth under arbitrary norms. We provide a streamlined convergence proof (Theorem 4.1) for our generalized algorithm and show that the number of oracle calls required for our binary search step is at most logarithmic. A key challenge in this setting is that our iterates may be unbounded, unlike in the Euclidean case. Additionally, we provide a lower bound for the binary search step when applied to smooth non-star-convex functions.
Nearly optimal performance for $p$-norm. For $p$-norms, we prove the same convergence rates as shown for convex functions by [NEMIROVSKII198521], extending them to star-convex functions. According to the lower bounds by [GUZMAN20151], our rates are nearly optimal, up to a logarithmic factor. Table 1 contextualizes our results in the existing literature.
@@
$(1,2]$ Upper $\mathcal{O}_{\kappa}\left(\big{(}\frac{LR^{\kappa}}{\epsilon}\big{)}^{\frac{2}{3\kappa-2}}\right)$ $\mathcal{O}_{\kappa}\left(\big{(}\frac{L\tau^{2}R^{\kappa}}{\epsilon}\big{)}^{\frac{2}{3\kappa-2}}\log^{2}\Big{(}\frac{L\tau R}{\epsilon}\Big{)}\right)$
[NEMIROVSKII198521] Our work (Corollary 5.2(i))
Lower $\Omega_{\kappa}\left(\big{(}\frac{LR^{\kappa}}{\epsilon[\ln d]^{\kappa-1}}\big{)}^{\frac{2}{3\kappa-2}}\right)$
[GUZMAN20151]
$[2,\infty]$ Upper $\mathcal{O}_{p,\kappa}\left(\big{(}\frac{LR^{\kappa}}{\epsilon}\big{)}^{\frac{p}{\kappa p+\kappa-p}}\right)$ $\mathcal{O}_{p,\kappa}\left(\big{(}\frac{L\tau^{p}R^{\kappa}}{\epsilon}\big{)}^{\frac{p}{\kappa p+\kappa-p}}\log^{2}\Big{(}\frac{L\tau R}{\epsilon}\Big{)}\right)$
[NEMIROVSKII198521] Our work (Corollary 5.2(ii))
Lower $\Omega_{p,\kappa}\left(\big{(}\frac{LR^{\kappa}}{\epsilon}\big{)}^{\frac{p}{\kappa p+\kappa-p}}\right)$
[GUZMAN20151]
Table 1: Summary of the relation between existing work and our results for $\tau$-star-convex functions. $R$ is the radius of the domain of the $\tau$-star-convex objective function, $(L,\kappa)$ is its smoothness with respect to a $p$-norm $\|\cdot\|_{p}$ with $p>1$ and $\epsilon$ the accuracy of the algorithm. Since convex functions are $1$-star-convex, the lower bounds from [GUZMAN20151] apply to both.
The article is organized as follows. In Section 2, we outline the problem setting and provide preliminaries. Section 3 presents our generalized accelerated algorithm and discusses its convergence guarantees for various settings. Section 4 introduces our binary search sub-algorithm and its running time. In Section 5 we put the two ingredients together and obtain our main bound. A discussion of future research directions is given in Section 6. All proofs are provided in the Appendix.
@@ 2.1 Definitions
We consider the class of $\tau$-star-convex functions, as introduced by [Hardt2016GradientDL], which are defined as follows.
@@
A continuously differentiable function $F:\mathbb{R}^{d}\xrightarrow{}\mathbb{R}$ is said to be $\tau$-star-convex for $\tau>0$ if there exists $x_{\star}\in\arg\min F$ such that for all $x\in\mathbb{R}^{d}$
$\tau\langle\nabla F(x),x-x_{\star}\rangle~{}\geq~{}F(x)-F(x_{\star}).$ (1)
Note that, all convex functions are $1$-star-convex [joulani17]. The condition becomes weaker as $\tau$ grows; conversely, if allowed, 00-star convex function would be identically constant. In our unconstrained optimization setting, we observe that $x_{\star}$ must be a global minimizer (as the gradient vanishes at global minima).
@@
For a continuously differentiable function $F:\mathbb{R}^{d}\xrightarrow{}\mathbb{R}$, the Bregman divergence from $x\in\mathbb{R}^{d}$ to $y\in\mathbb{R}^{d}$ is defined by
@@
Note that we do not assume convexity of $F$, and consequently we may have $D_{F}(x,y)<0$.
@@
The special case $\kappa=2$ is simply referred to as smooth.
To develop our mirror descent algorithm, we use a so-called distance-generating function (see Assumption 1), which fulfils a strong or uniform convexity assumption with respect to an arbitrary norm, defined as follows.
@@
The special case $q=2$ is called strongly convex.
The classic acceleration setting considers $q=\kappa=2$. In general, we have $q>2>\kappa$.
@@ 3 Main results
Our accelerated optimization algorithm for weakly smooth, star-convex functions incorporates both a mirror descent step and a proximal step, similar to [AGJ18]. However, as we will see in the proof sketch, our analysis exploits the specific structure of unconstrained optimization. We start by summarizing our assumptions.
@@
For an arbitrary norm, constructing a distance-generating function that satisfies the uniform convexity assumption of Definition 2.4 requires careful consideration. Below we present examples of distance-generating functions for $p$-norms and composite norms, each meeting the required conditions of uniform convexity.
Example 3.1 ($p$-norms).
To construct the corresponding distance-generating function, we define $\psi(x)=\frac{1}{q}\|x\|_{p}^{q}$, where the appropriate order $q$ depends on $p$. Specifically, for $p\in(1,2]$, we set $q=2$, ensuring that $\psi$ is $(p-1)$-strongly convex with respect to $\left\lVert\cdot\right\rVert_{p}$ [uniformconvex1, uniformconvex2]. For $p>2$, we choose $q=p$, ensuring that $\psi$ is $(2^{-\frac{p(p-2)}{p-1}},p)$-uniformly convex with respect to  $\|\cdot\|_{p}$ [[, see]]Zalinescu:1983,AGL24.
For both cases, we will show sharp convergence guarantees of our algorithm, as summarized in Table 1.
Example 3.2 (Composite norms and distance generating function).
For composite norms, the distance-generating function is constructed by combining the distance-generating functions associated with the individual norms, ensuring the desired uniform convexity property. Consider, for instance, the following composite norm
$\|x\|_{2\circ 1.5}~{}:=~{}\sqrt{\frac{1}{2}\|x_{1:d/2}\|^{2}_{2}+\frac{1}{2}\|x_{(d/2+1):d}\|^{2}_{1.5}}.$ (2)
By Example 3.1, $\psi_{1}(x_{1:d/2})=\frac{1}{2}\sum_{i=1}^{d/2}x_{i}^{2}$ is $1$-strongly convex with respect to $\|\cdot\|_{2}$ and $\psi_{2}(x_{d/2+1,d})=\frac{1}{1.5}(\sum_{i=d/2+1}^{d}|x_{i}|^{1.5})^{2/1.5}$ is $\frac{1}{2}$-strongly convex with respect to $\|\cdot\|_{1.5}$. Hence, we define the composite distance-generating function as $\psi(x):=\frac{1}{2}\psi_{1}(x_{1:d/2})+\frac{1}{2}\psi_{2}(x_{d/2+1,d}).$ For any $x,y\in\mathbb{R}^{d}$, we denote the Bregman divergence by $D_{\psi}(x,y)$ and we note that
$D_{\psi}(x,y)$ $=\frac{1}{2}\Big{(}\psi_{1}(x_{1:d/2})-\psi_{1}(y_{1:d/2})-\langle\nabla\psi_{1}(y_{1:d/2}),x_{1:d/2}-y_{1:d/2}\rangle\Big{)}$
$\quad+\frac{1}{2}\Big{(}\psi_{2}(x_{(d/2+1):d})-\psi_{2}(y_{(d/2+1):d})-\langle\nabla\psi_{2}(y_{(d/2+1):d}),x_{(d/2+1):d}-y_{(d/2+1):d}\rangle\Big{)}$
$\geq\frac{1}{4}\|x_{1:d/2}-y_{1:d/2}\|_{2}^{2}+\frac{1}{8}\|x_{(d/2+1):d}-y_{(d/2+1):d}\|_{1.5}^{2}~{}\geq~{}\frac{\|x-y\|_{2\circ 1.5}^{2}}{4}.$
We use the strong convexity of $\psi_{1}$ and $\psi_{2}$ in the last line. Thus, $\psi$ is $\frac{1}{2}$-strongly convex with respect to the composite norm $\|\|_{2\circ 1.5}$. In general, the strong convexity coefficient of the composite norm is determined by the weaker of the two individual norms.